OpenStudy (anonymous):

need help with this 2/9k+(-1/5)=2-(-2/7k), solve for k

OpenStudy (anonymous):

\[\frac{ 2 }{ 9 }k+\left( \frac{ -1 }{ 5 } \right)=2-\left( \frac{ -2 }{ 7 } \right)k\]

OpenStudy (anonymous):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

(2/9)k-(2/7)k = 2+(1/5) k((2/9)-(2/7)) = 11/5 k(-4/63) = 11/5 k = (11/5)(63/-4) = -693/20